# AIJack: Security and Privacy Risk Simulator for Machine Learning

### 29 Dec 2023 v1

@@ Abstract
This paper introduces AIJack, an open-source library designed to assess security and privacy risks associated with the training and deployment of machine learning models. Amid the growing interest in big data and AI, advancements in machine learning research and business are accelerating. However, recent studies reveal potential threats, such as the theft of training data and the manipulation of models by malicious attackers. Therefore, a comprehensive understanding of machine learning’s security and privacy vulnerabilities is crucial for the safe integration of machine learning into real-world products. AIJack aims to address this need by providing a library with various attack and defense methods through a unified API. The library is publicly available on GitHub (https://github.com/Koukyosyumei/AIJack).
@@ 1 Introduction
Machine learning (ML) has become a foundational component of diverse applications, spanning image recognition to natural language processing. As these technologies proliferate, the need to comprehend and address security risks associated with ML models becomes imperative.
While ML models continuously enhance accuracy, attackers can significantly diminish it by introducing artificially created data. Evasion Attacks [1, 2], using specialized algorithms, can induce malfunctions in models. Consider a road sign classification model; by adding imperceptible noise to a ”speed limit 30km” sign, it may misclassify it as ”speed limit 60km.” Another attack, Poisoning Attack [3], injects contaminated data during training, lowering model accuracy.
To counter such attacks, various strategies have been proposed. Certified robustness [4, 5], a technique to formally guarantee that adversarial examples cannot lead to undesirable predictions, has gained attention. Another approach is debugging machine learning models [6, 7, 8], which aims to identify inputs causing unexpected behaviors.
@@
Privacy protection techniques, including differential privacy [12], k-anonymity [13], homomorphic encryption [14], and distributed learning [15, 16], have been proposed. Differential privacy prevents individual data inference, while homomorphic encryption enables arithmetic operations on encrypted data. Distributed methods like Federated Learning (FL) [15] and Split Learning [16] facilitate collaborative learning among data owners.
It is crucial to assess the security and privacy risks of ML models and evaluate countermeasure effectiveness. To simplify such simulations, we propose the open-source software, AIJack, offering various attack and defense methods. AIJack enables experimenting with various combinations of attacks and defenses with simple code. Built on PyTorch [17] and scikit-learn [18], users can easily incorporate AIJack into existing code.
@@ 2 Package Design
Collaborative Horizontal FL FedAVG [15], FedProx [19], FedMD [20], FedGEMS [21], DSFL [22], MOON [23], FedEXP [24]
Collaborative Vertical FL SplitNN [16], SecureBoost [25]
Attack Model Inversion MI-FACE [9], DLG [10], iDLG [26], GS [27], CPL [28], GradInversion [29], GAN Attack [30]
@@
Attack Poisoning History Attack [32], Label Flip [32], MAPF [32], SVM Poisoning [3]
@@
Defense Robust Training PixelDP [4], Cost-Aware Robust Tree Ensemble [5]
@@
Code 1 and Code 2 show the example codes implementing Evasion Attack and Poisoning Attack against SVM trained with scikit-learn, respectively.
1from aijack.attack import (
2    Evasion_attack_sklearn
3)
@@ 2    Evasion_attack_sklearn
5attacker = Evasion_attack_sklearn(
6    clf, initial_data_positive
7)
8malicious_data, log = attacker.attack(
9    initial_data_negative
10)
Code 1: Evasion Attack against SVM
1from aijack.attack import (
2    Poison_attack_sklearn
3)
4
5attacker = Poison_attack_sklearn(
6    clf, X_train, y_train
7)
8malicious_data, log = attacker.attack(
9    initial_data, 1, X_valid, y_valid
10)
Code 2: Poisoning Attack against SVM
Example codes for Federated Learning can be found in Code 3 to Code 6. Code 3 implements FedAVG with PyTorch, the standard Federated Learning protocol, Code 4 implements a Gradient Inversion Attack against Federated Learning (a type of Model Inversion Attack), Code 5 applies Paillier Encryption, one of the most popular Homomorphic Encryption methods, to mitigate inversion attacks, and Code 6 migrates Code 3 to MPI-backend.
@@ 4  FedAVGAPI
15clients = [FedAVGClient(model_1),
16           FedAVGClient(model_2)]
17server = FedAVGServer(clients, model_3)
18
19api = FedAVGAPI(
@@ 16           FedAVGClient(model_2)]
22)
23api.run()
Code 3: Standard Federated Learning (FL)
@@ 7  GradientInversionAttackServerManager
12)
13MaliciousServer = mg.attach(FedAVGServer)
14
15clients = [FedAVGClient(model_1),
16           FedAVGClient(model_2)]
17server = MaliciousServer(clients, model_3)
@@ 16           FedAVGClient(model_2)]
19api = FedAVGAPI(
@@
22)
23api.run()
Code 4: Gradient Inversion Attack against FL
@@ 8  PaillierKeyGenerator
17clients = [PaillierClient(model_1),
18           PaillierClient(model_2)]
19server = FedAVGServer(clients, model_3)
@@ 18           PaillierClient(model_2)]
Code 5: FL with Paillier Encryption
1from aijack.collaborative import (
2  FedAVGClient,
3  FedAVGServer,
4  MPIFedAVGAPI,
5  MPIFedAVGClientManager,
6  MPIFedAVGServerManager
@@
8
9mcm = MPIFedAVGClientManager()
10msm = MPIFedAVGServerManager()
11MPIClient = mcm.attach(FedAVGClient)
12MPIGServer = msm.attach(FedAVGServer)
@@
14if myid == 0:
15    client_ids = list(range(1, size))
16    server = MPIServer(comm, client_ids, model)
17    api = MPIFedAVGAPI(
18        comm, server, True, criterion
19    )
20else:
21    client = MPIClient(comm, model, user_id=myid)
22    api = MPIFedAVGAPI(
23        comm, client, False, criterion,
24        optimizer, dataloader
25    )
26
27api.run()
Code 6: FL with MPI backend
Code7 illustrates the standard workload involved in training a neural network using Differential Private Stochastic Gradient Descent (DPSGD). Our API’s design draws inspiration from Opacus [43], a renowned framework dedicated to deep learning with differential privacy. Additionally, our AIJack facilitates a more intuitive implementation, closely aligning with the algorithm initially proposed in[12].
@@
5
6accountant.reset_step_info()
7dpoptimizer_cls, lot_loader, batch_loader = privacy_manager.privatize(
@@
9)
10optimizer = dpoptimizer_cls(net.parameters(), lr=lr)
11
12for epoch in range(num_iterations):
@@
16            loss = criterion(net(X_batch), y_batch)

### 08 Apr 2024 v2

# AIJack: Let’s Hijack AI! Security and Privacy Risk Simulator for Machine Learning

@@ Abstract
This paper introduces AIJack, an open-source library designed to assess security and privacy risks associated with the training and deployment of machine learning models. Amid the growing interest in big data and AI, machine learning research and business advancements are accelerating. However, recent studies reveal potential threats, such as the theft of training data and the manipulation of models by malicious attackers. Therefore, a comprehensive understanding of machine learning’s security and privacy vulnerabilities is crucial for the safe integration of machine learning into real-world products. AIJack aims to address this need by providing a library with various attack and defense methods through a unified API. The library is publicly available on GitHub (https://github.com/Koukyosyumei/AIJack).
@@ 1 Introduction
Machine learning (ML) has become a foundational component of diverse applications, spanning image recognition to natural language processing. As these technologies proliferate, comprehending and addressing security risks associated with ML models becomes imperative.
While ML models continuously enhance accuracy, attackers can significantly diminish it by introducing artificially created data. Evasion Attacks [1, 2], using specialized algorithms, can induce model malfunctions. Consider a road sign classification model; adding imperceptible noise to a ”speed limit 30km” sign may misclassify it as ”speed limit 60km.” Those perturbed samples are often called adversarial examples. Another attack, Poisoning Attack [3], injects contaminated data during training, lowering model accuracy.
Various strategies have been proposed to counter such attacks. Certified robustness [4, 5], a technique to formally guarantee that adversarial examples cannot lead to undesirable predictions, has gained attention. Another approach is debugging machine learning models [6, 7, 8], which aims to identify inputs causing unexpected behaviors.
@@
Privacy protection techniques, including differential privacy [12], k-anonymity [13], homomorphic encryption [14], and distributed learning [15, 16], have been proposed. Differential privacy prevents individual data inference, while homomorphic encryption enables arithmetic operations on encrypted data. Distributed methods like federated learning (FL) [15] facilitate collaborative learning among data owners without violating data privacy.
Thus, assessing ML models’ security and privacy risks and evaluating countermeasure effectiveness is crucial. To simplify such simulations, we propose the open-source software, AIJack, offering various attack and defense methods. AIJack enables the experimentation of numerous combinations of attacks and defenses with simple code. Built on PyTorch [17] and scikit-learn [18], users can easily incorporate AIJack into existing code.
@@ 2 Package Design
FL Horizontal
FedAVG [15], FedProx [19], FedMD [20],
FedGEMS [21], DSFL [22], MOON [23], FedEXP [24]
FL Vertical SplitNN [16], SecureBoost [25]
Attack Model Inversion
MI-FACE [9], DLG [10], iDLG [26], GS [27],
CPL [28], GradInversion [29], GAN Attack [30]
@@
Attack Poisoning
History Attack [32], Label Flip [32],
MAPF [32], SVM Poisoning [3]
@@
Defense Robustness PixelDP [4], Cost-Aware Robust Tree Ensemble [5]
@@
Examples
For example, users can easily experiment with evasion attacks and mitigation methods. Code 1 shows FGSM, one of the most popular evasion attacks against a neural network, while Code 2 demonstrates PixelDP, one of the most popular mitigation methods against evasion attacks providing certified robustness. These codes can be integrated on the top of PyTorch-based machine learning pipelines.
1from aijack.attack import FGSMAttacker
2
3
@@
5attacker = FGSMAttacker($n$, criterion, eps=0.3)
6perturbed_x = attacker.attack((x, y))
Code 1: FGSM attack against a neural network
1from aijack.defense import PixelDP
2
3pdp = PixelDP(NN(), num_classes)
4pdp.eval()
5pred_proba, counts = pdp(data)
6pdp.certify(counts)
Code 2: PixelDP
Code. 3 shows the example code implementing the MIFACE attack, one of the most popular model inversion attack methods based on gradient descent, allowing attackers to steal training data from the trained model.
1from aijack.attack import MI_FACE
2
3attacker = MI_FACE($n$, input_shape)
4recovered_data, log = attacker.attack()
Code 3: MIFACE against NN
One possible defense method against such model inversion attacks is training a neural network using Differential Private Stochastic Gradient Descent (DPSGD). DPSGD is easily implemented with AIJack on the top of PyTorch, as illustrated in Code 4.
2
4privacy_manager = DPSGDManager(accountant, optim.SGD)
5
6accountant.reset_step_info()
7dpoptimizer, lot_loader, batch_loader = privacy_manager.privatize(noise_multiplier=sigma)
8optimizer = dpoptimizer($n$.parameters(), lr=lr)
9
10for epoch in range(num_iterations):
11    for X_lot, y_lot in lot_loader(optimizer):
13            optimizer.zero_grad()
14            loss = criterion(nn(X_batch), y_batch)
15            loss.backward()
16            optimizer.step()
Code 4: Deep Learning with Differential Privacy
AIjack also makes it easy to implement various types of federated learning schemes and simulate their security and privacy risks. Code 5 shows the implementation of FedAVG, one of the most standard federated learning protocols, on top of PyTorch. Code 6 mitigates Code 5 to MPI-backend.
@@ 4  FedAVGAPI
15
16
17clients = [FedAVGClient(NN()),
18           FedAVGClient(NN())]
19server = FedAVGServer(clients, NN())
20
21api = FedAVGAPI(
@@ 18           FedAVGClient(NN())]
24)
25api.run()
Code 5: Standard FedAVG
1from aijack.collaborative import (
2  FedAVGClient, FedAVGServer, MPIFedAVGAPI,
3  MPIFedAVGClientManager,
4  MPIFedAVGServerManager
5)
6
7mcm = MPIFedAVGClientManager()
8msm = MPIFedAVGServerManager()
9MPIClient = mcm.attach(FedAVGClient)
10MPIGServer = msm.attach(FedAVGServer)
11
12if myid == 0:
13    ids = list(range(1, size))
14    server = MPIServer(com, ids, NN())
15    api = MPIFedAVGAPI(
16        comm, server, True, criterion
17    )
18else:
19    client = MPIClient(comm, NN(), myid)
20    api = MPIFedAVGAPI(
21        comm, client, False, criterion,
22        optimizer, dataloader
23    )
24
25api.run()
Code 6: FedAVG with MPI backend
Then, users can easily incorporate various attack and defense strategies into federated learning by attaching those abilities to servers and clients. For instance, Code 7 demonstrates a gradient-based model inversion attack against federated learning, while Code 8 applies Paillier Encryption, one of the most popular Homomorphic Encryption methods, to mitigate such inversion attacks.
@@ 7  GradientInversionAttackServerManager
16)
17MaliciousServer = mg.attach(FedAVGServer)
@@
19clients = [FedAVGClient(NN()),
20           FedAVGClient(NN())]
21server = MaliciousServer(clients, NN())
22
23api = FedAVGAPI(
@@ 20           FedAVGClient(NN())]
26)
27api.run()
Code 7: Gradient Inversion Attack against FL
@@ 8  PaillierKeyGenerator
17clients = [PaillierClient(NN()),
18           PaillierClient(NN())]
19server = FedAVGServer(clients, NN())
@@ 18           PaillierClient(NN())]
Code 8: FL with Paillier Encryption
Additional examples are available in the Appendix and our documentation [^1] https://koukyosyumei.github.io/AIJack/.
Comparison with other libraries
While many existing libraries prioritize security and privacy within the realm of machine learning, AIJack distinguishes itself by offering unparalleled flexibility and a wide array of options for exploring security and privacy attacks and their corresponding countermeasures. Unlike many prior tools that focus on specific types of attacks and defenses—such as ART [43], foolbox [44], AdvBox [45], and advertorch [46] for adversarial examples; ml_privacy_meter [47] and PrivacyRaven [48] for model inversion and membership inference attacks; tensorflow-privacy [49] and Opacus [50] for differential privacy; and FATE [51], flower [52], FedML [53], VFLAIR [54], and secretflow [55] for federated learning—AIJack stands out by encompassing all of these and more. This comprehensive approach empowers users to simulate various attack scenarios easily and experiment with multiple defense methods simultaneously.
Acknowledgement
We acknowledge every contributor for their support of this project. The comprehensive list of contributors is in our GitHub repository.
Appendix A Standard Machine Learning
This section explores the vulnerabilities and defenses of standard machine learning. We delve into various attack techniques, including adversarial examples, poisoning attacks, model inversion attacks, and membership inference attacks. We then discuss defense mechanisms like certified robustness, debugging methods, differential privacy, and k-anonymity, all crucial for building secure and trustworthy machine learning systems.
Furthermore, practical examples and code snippets are provided to illustrate the implementation of both attack and defense strategies with AIJack. Since some detailed arguments and parameters are committed for better readability, it might be better for interested readers to check the detailed documentation at https://koukyosyumei.github.io/AIJack/.
A.1 Attack
Adversarial Examples
Adversarial examples are carefully crafted inputs aimed at deceiving machine learning models [56]. These subtle alterations, often undetectable to humans, can lead the model to produce noticeably incorrect outputs. This form of attack is also known as an evasion attack. For example, an adversarial example could be a slightly altered image that causes a computer vision system to misidentify it as an entirely different object. Consequently, adversarial examples raise concerns about the reliability and applicability of machine learning models in real-world scenarios. Ongoing research in this domain focuses on techniques for generating and defending against adversarial examples, promoting the development of more resilient and dependable machine learning systems.
Code S-1 shows the example codes implementing the gradient-based evasion attack [1] against SVM trained with scikit-learn, and Code S-2 demonstrates the FGSM attack [2] against a neural network implemented with PyTorch.
1from aijack.attack import (
2    Evasion_attack_sklearn
3)
4
5attacker = Evasion_attack_sklearn(
6    clf, initial_data_positive
@@ 2    Evasion_attack_sklearn
8malicious_data, log = attacker.attack(
9    initial_data_negative
10)
Code S-1: Evasion Attack against SVM
1from aijack.attack.evasion import FGSMAttacker
2
3
4
5attacker = FGSMAttacker(
6    $n$, criterion, eps=0.3
7)
8perturbed_x = attacker.attack((x_origin, y_origin))
Code S-2: FGSM attack against NN
Poisoning Attack
In machine learning, poisoning attacks are a significant threat aimed at manipulating a model’s training process toward malicious ends [57]. They often involve injecting corrupted data points, targeting specific classes to reduce accuracy for certain predictions, or even altering the model’s decision boundaries. For example, injecting mislabeled images into a facial recognition model could lead to the misidentification of specific individuals.
Code S-3 is the example code of the Poisoning Attack against SVM [3] trained with scikit-learn.
1from aijack.attack import Poison_attack_sklearn
2
3attacker = Poison_attack_sklearn(clf, X_train, y_train)
4malicious_data, log = attacker.attack(initial_data, 1, X_valid, y_valid)
Code S-3: Poisoning Attack against SVM
Model Inversion Attack
Model inversion attacks present a serious security risk in machine learning, especially with sensitive data. They involve reconstructing training data by analyzing a model’s parameters and outputs, often using techniques like gradient descent on carefully crafted queries [58]. Successful attacks can compromise sensitive information in the training data, risking user privacy or disclosing confidential details.
Code. S-4 shows the example code implementing the MIFACE attack [9], one of the most popular model inversion attack methods based on gradient descent.
1from aijack.attack import MI_FACE
2
3mi = MI_FACE($n$, input_shape)
4x_result_1, log = mi.attack()
Code S-4: MIFACE against NN
Membership Inference Attack
Machine learning models are also vulnerable to membership inference attacks (MIAs), where attackers try to ascertain if a data point is part of the model’s training dataset [59]. By analyzing the model’s outputs, like predictions or confidence scores, attackers can craft queries to discern membership. Deviations in the model’s behavior for a specific data point compared to unseen data may reveal membership in the training set. MIAs are a significant privacy concern, particularly with sensitive datasets such as medical records or financial transactions.
Code. S-5 shows the example implementation of a shadow attack [11], which is one of the earliest and most popular membership inference attacks.
1from aijack.attack.membership import ShadowMembershipInferenceAttack
2
3shadow_models = [clf() for _ in range(2)]
4attack_models = [clf() for _ in range(10)]
5
6attacker = ShadowMembershipInferenceAttack(clf, shadow_models, attack_models)
7attacker.fit(X_shadow, y_shadow)
Code S-5: Membership Inference Attack
A.2 Defense
Certified Robustness
Certified robustness measures a model’s assured resilience against adversarial examples, offering formal guarantees beyond traditional accuracy metrics [60, 61]. To achieve this, methods like randomized smoothing inject noise into the internal process, though it may lower accuracy and computational efficiency. Ongoing research aims to enhance certification methods, balancing robustness, accuracy, and efficiency for practical machine-learning applications.
Code S-6 is the example of PixelDP [4] applied for a neural network implemented with PyTorch. Code S-7 also shows Cost-Aware Robust Tree Ensemble [5], which is the tree-based certified robust training method.
1from aijack.defense import (
2    PixelDP
3)
4
5pdp = PixelDP(
6    $n$,
7    num_classes,
8    eps=0.3,
9    delta=1e-5,
10)
11
12# Normal Training of PyTorch
@@ 6    $n$,
14pdp.eval()
15pred_proba, counts = pdp(data)
16pdp.certify(counts)
Code S-6: PixelDP
1from aijack.collaborative.tree import (
2    XGBoostClassifierAPI,
3    XGBoostClient,
4)
5
6
7
8p0 = XGBoostClient(X_train)
9
10# Set the attack-cost constraint to each feature
11p0.set_cost_constraint_map(constraints)
12
13clf = XGBoostClassifierAPI()
14clf.fit([p0], y_train.tolist())
15
16pred_proba = clf.predict_proba(X_test)
Code S-7: Cost-Aware Robust Tree Ensemble
Debugging
Debugging neural networks is uniquely challenging due to their complex, non-linear nature. Neuron coverage [8], akin to code coverage in software engineering, measures the proportion of activated neurons above a predefined threshold during testing. Higher coverage reflects a more thorough exploration of the network’s decision space. This method can be utilized to detect unknown adversarial examples and unexpected behaviors of a neural network caused by poisoning attacks.
With AIJack, users can easily implement Neuron Coverage, as shown in Code S-8.
1from aijack.defense.debugging.neuroncoverage import *
2
3NCT = neuroncoverage.NeuronCoverageTracker($n$, threshold, dummy_data)
4nc = NCT.coverage(debug_dataloader)
Code S-8: Debugging of a neural network with Neuron Coverage
Differential Privacy
Differential privacy (DP) offers a formal privacy guarantee in randomized algorithms [62]. It ensures that the inclusion or exclusion of a single data point has minimal impact on the algorithm’s output. Applying differential privacy to machine learning protects individual privacy while still enabling the training of effective models [63]. This method can be the solution against the model inversion attack and membership inference attacks.
Differentially, Private Stochastic Gradient Descent (DPSGD) [12] is a prominent technique that injects noise into the gradient updates during model training. This added noise safeguards individual data points while introducing a controllable privacy-utility trade-off. DPSGD serves as a cornerstone for privacy-preserving machine learning, allowing researchers to train models on sensitive data without compromising individual privacy.
Code S-9 illustrates the standard workload of training a neural network using Differential Private Stochastic Gradient Descent (DPSGD). AIJack also supports several variants of DPSGD, including AdaDPS [37] and DPlis [38]. Our API’s design draws inspiration from Opacus [50], a renowned framework dedicated to deep learning with differential privacy. Additionally, our AIJack facilitates a more intuitive implementation, closely aligning with the algorithm initially proposed in [12].
@@ 3    XGBoostClient,
5# privacy_manager = AdaDPSManager(accountant, optim.SGD)
6
7accountant.reset_step_info()
8dpoptimizer_cls, lot_loader, batch_loader = privacy_manager.privatize(
@@
10)
11optimizer = dpoptimizer_cls($n$.parameters(), lr=lr)
12
13for epoch in range(num_iterations):
@@
17            loss = criterion(nn(X_batch), y_batch)